OpenStudy (anonymous):

the radioactive substance krypton-91 has a very short half life of about 10 seconds what is the decay constant for this substance

OpenStudy (psymon):

k = -ln2/(half-life). Quick formula for those exponential growth and decay problems.

OpenStudy (anonymous):

\[N(t) = No \frac{ 1 }{ 2 }^{-\lambda t}\] Is this the formula you use ?

OpenStudy (anonymous):

\[N(t) = N_0 e^{- \lambda t}\]

OpenStudy (unklerhaukus):

\[\lambda=\frac{\ln(2)}{t_{1/2}}\]

OpenStudy (anonymous):

ok thanks UnkleRhaukus. Ill use that one since im not really sure what formula to use since the problem doesnt tell me what formula to use.